ES014B

PI: Shen
Title: 0106+013

Session: Feb 1997 (session 1)
Frequency: 6 cm
Mode: A/PI
Obs time (UT): EU 09:00(25/02)-16:00(25/02)
Stations: EbHhJbMcNtOnShTrUrWb
Processor: Bonn
EU Mk3 Tapes/St.: 2

Comment


To give feedback on this experiment, click here

Experiment Feedback

Onsala (Antonis Polatidis at Wed Feb 26 08:54:57 EST 1997) Failure
Due to high winds (up to 20m/s) we could not observe at all during this experiment. We made two attempts to unstow the antenna but the tracking was awful

Westerbork (Tony Foley at Fri Feb 28 08:52:33 EST 1997) Minor failures
Used 8 telescopes. Shadowing of telescopes for the scans 0908 to 0945UT. Scan scheduled for 1053 UT out of hourangle range. Shadowing of telescopes 1230-1235UT. Tape change at 1241 was 3.5min late. One telescope (RT8) offsource around 1345

HartRAO (Jonathan Quick at Sat Mar 1 17:24:42 EST 1997) Minor failures
Read heads failing - Write head status unknown Lost one scan due to FS9 crash

Effelsberg (dgraham at Mon Mar 3 03:17:59 EST 1997) Success
OK

Shanghai (Huang Xinyong at Wed Mar 5 05:13:16 EST 1997) Success
Clock offset = 0.39-0.40 msec. No problems, weather overcast.

Jodrell (Alastair Gunn at Thu Mar 6 04:10:51 EST 1997) Failure
Did not observe due to high winds throughout experiment.

Calibration (Fredrik T Rantakyro at Thu May 8 11:33:01 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the Feb97 directory with the name es14b.antab. This file contains the preliminary calibration results and may be updated when I receive updated information from the telescopes and the results from the datareduction of the calibration experiments.

MPIfR (Scott Aaron at Fri Mar 27 11:06:43 MET 1998) Minor failures
Correlation completed. No fringes found to HartRAO. No other problems found.